David is two years older than peter and five years older than John.Their ages add up to 44. How old is John?
Found 3 solutions by Theo, ikleyn, MathTherapy:
Answer by Theo(13342)

let d = david's age.
let p = peter's age.
let j = john's age.

david is 2 years older than peter, therefore d = p + 2.
david is 5 years older than john, therefore d = j + 5.

you are given that d + p + j = 44

in d = p + 2, solve for p to get:
p = d - 2.

in d = j + 5, solve for j to get:
j = d - 5.

replace p and j with their equivalent values in d to get:
d + p + j = 44 becomes:
d + (d - 2) + (d - 5) = 44
combine like terms to get:
3 * d - 7 = 44
add 7 to both sides of the equation to get 3 * d = 51
solve for d to get d = 51 / 3 = 17.

since d = p + 2, then p = 15.
since d = j + 5, then j = 12.

d + p + j becomes 17 + 15 + 12 = 44

the values for d and p and j are confirmed to be good.

the solution is that john is 12 years old.


Answer by ikleyn(52788)

.
Let x be the David's age.


Then Peter is (x-2) years old, while John is (x-5) years old.


The total of their ages is 44, which gives you an equation


    x + (x-2) + (x-5) = 44,

Or

    3x - 7 = 44,

    3x = 44 + 7 = 51,

     x           = 51/3 = 17.


ANSWER.  John's age is  x-5 = 17-5 = 12 years.

This problem deserves to be solved in short,  simple and clear form.

As any other problem,  too . . .



Answer by MathTherapy(10552)

David is two years older than peter and five years older than John.Their ages add up to 44. How old is John?
Let John's age be J
Then David's is: J + 5, and Paul's is: J + 5 - 2, or J + 3
With ages totaling 44 years, we get: J + J + 5 + J + 3 = 44
3J = 36
J, or